Solve using the quadratic formula
2x squared -3x -5=0

Answer by stanbon(75887)

Solve using the quadratic formula
2x squared -3x -5=0
--------------
2x^2 - 3x - 5 = 0
-----
x = [-b +-sqrt(b^2-4ac)]/(2a)
-----
x = [3 +- sqrt(9-4*2*-5)]/(4)
-----
x = [3 +- sqrt(49)]/4
-----
x = [3+7]/4 = 5/2
OR
x = [3-7]/4 = -1
------------
Cheers,
Stan H.
=============